2.A pool company forgets to bring their logarithmic charts, but they need to raise the amount of hydronium ions in a pool by 0.50. Using complete sentences, explain how your graph can be used to solve 10–y = 0.50. Find the approximate solution.
Answer by stanbon(75887)

.A pool company forgets to bring their logarithmic charts, but they need to raise the amount of hydrogen ions in a pool by 0.50. Using complete sentences, explain how your graph can be used to solve 10^(–y) = 0.50. Find the approximate solution
===========
Take the log of both sides to get::
-y = log(0.5)
---
y = -log(0.5)
----
y = -(-0.3010)
----
y = 0.3010
---------------
Cheers,
Stan H.